# Waveform Design of Multi-User-Multi-Target ISAC System based on Kullback–Leibler Divergence

### v1

@@ VII-C Communication Waveform KLD-based Optimisation
The communication waveform KLD-based optimization algorithm is configured the same as the radar one with a maximum iteration limit of $\mathrm{max\_iter}=1000$ and a convergence tolerance of $\epsilon=10^{-6}$. The barrier parameter is initialized at $\tilde{\mu_{0}}=1$, with a barrier reduction factor of $\tilde{\gamma}=0.1$. The KLD lower bounds $\underline{A}_{t}$ and $\underline{B}_{k}$ for all targets and UEs are set to 10 bits.
Fig. 8 illustrates the performance of the radar and communication subsystems when using the optimized communication waveform. Fig. 8.a shows the achievable $\mathrm{KLD_{r}}$ for each target as a function of $P_{\mathrm{R}}/N_{o}$ for three different power allocation cases: $P_{\mathrm{r}}=\{0.7,0.5,0.3\}$. As $P_{\mathrm{R}}/N_{o}$ increases, $\mathrm{KLD_{r}}$ improves across all targets in each power allocation scenario, demonstrating enhanced radar performance at higher power-to-noise ratios. For each power allocation, target $\mathrm{T}_{3}$ consistently achieves the highest $\mathrm{KLD_{r}}$, followed by $\mathrm{T}_{2}$ and $\mathrm{T}_{1}$, indicating that the radar performance is influenced by their respective distances or scattering characteristics. Specifically, at $P_{\mathrm{R}}/N_{o}=10$ dB and $P_{\mathrm{r}}=0.7$, the $\{\mathrm{KLD}_{\mathrm{r},1},\mathrm{KLD}_{\mathrm{r},2},\mathrm{KLD}_{\mathrm{r},3}\}$ values are approximately $\{8.42,10.78,12.64\}$ for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$, respectively. These values increase to about $\{38.26,42.14,47.53\}$ at $P_{\mathrm{R}}/N_{o}=20$ dB. The linearity observed in $\mathrm{KLD_{r}}$ improvement beyond 15 dB is consistent across the different $P_{\mathrm{r}}$ values, suggesting stable radar system performance at higher $P_{\mathrm{R}}/N_{o}$ values, despite the suboptimal radar waveform.
Fig. 8.b depicts the communication subsystem performance as represented by $\mathrm{KLD_{c}}$ versus $P_{\mathrm{R}}/N_{o}$ for three different power allocation cases: $P_{\mathrm{c}}=\{0.7,0.5,0.3\}$. $\mathrm{UE}_{3}$ consistently achieves the highest $\mathrm{KLD_{c}}$, followed by $\mathrm{UE}_{2}$ and $\mathrm{UE}_{1}$, within each power allocation scenario. For $P_{\mathrm{R}}/N_{o}=10$ dB and $P_{\mathrm{c}}=0.7$, the $\{\mathrm{KLD}_{\mathrm{c},1},\mathrm{KLD}_{\mathrm{c},2},\mathrm{KLD}_{\mathrm{c},3}\}$ values are approximately $\{45.87,63.92,86.14\}$ for $\mathrm{UE}_{1}$, $\mathrm{UE}_{2}$, and $\mathrm{UE}_{3}$, respectively. These values increase to around $\{128.43,159.47,182.76\}$ at $P_{\mathrm{R}}/N_{o}=20$ dB. The communication performance shows less sharp improvement at higher $P_{\mathrm{R}}/N_{o}$ due to the ISAC system’s effort to balance the radar subsystem thresholds with communication subsystem enhancements across different power allocation scenarios.
@@
Fig. 9.a shows the ROC curves for the three targets across different $P_{\mathrm{R}}/{N}o$ scenarios. The ROC curves plot the probability of detection ($P_{D}$) versus the probability of false alarm ($P_{FA}$) for each target.At $P_{\mathrm{R}}/{N}_{o}=10$ dB, and $P_{FA}=10^{-3}$, the $P_{D}$ is $\{0.397,0.231,0.454\}$ for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$, respectively, while for the non-optimal way the $P_{D}$ is $\{0.217,0.109,0.273\}$. The results demonstrate that communication optimization slightly improves detection capabilities at lower $P_{\mathrm{R}}/{N}_{o}$ compared to non-optimization, as seen in Fig. 3.
Fig. 9.b depicts the BER performance for three UEs versus $P_{\mathrm{R}}/N_{o}$. The BER decreases as $P_{\mathrm{R}}/N_{o}$ increases for all three UEs, indicating improved communication performance at higher $P_{\mathrm{R}}/N_{o}$. $\mathrm{UE}_{3}$ consistently achieves the lowest BER, followed by $\mathrm{UE}_{2}$ and $\mathrm{UE}_{1}$. The performance of the communication subsystem improves significantly with optimized communication waveform. In the non-optimal case, $P_{e}$ saturates at $8\times 10^{-3}$ after 15 dB, as shown in Fig. 2. However, with optimization, $P_{e}$ drops below $10^{-4}$ by $15$ dB across all UEs, optimized waveforms still provides a gain in comparison to the non-optimal method, however the gain is lower than optimised radar waveform due to the radar interference.
@@
Fig. 10.a illustrates the trade-off between radar and communication subsystems using the optimized communication waveform. The plot shows $\mathrm{KLD_{c}}$ versus $\mathrm{KLD_{r}}$, with each point representing a different $P_{\mathrm{R}}/N_{o}$ value. Moving from left to right, both $\mathrm{KLD_{c}}$ and $\mathrm{KLD_{r}}$ increase, indicating that higher $P_{\mathrm{R}}/N_{o}$ values enhance performance in both subsystems. This visualization highlights the robustness of the optimization technique in balancing radar and communication performance across different $P_{\mathrm{R}}/N_{o}$ levels, we can observe that this optimal way shows an increasing and favourable trade-off, by comparison to the non optimal way in Fig. 4.b, where the $\mathrm{KLD_{c}}$ saturates at around $11$ bits as $P_{\mathrm{R}}/N_{o}$ increases.
Fig. 10.b shows the performance of the radar waveform optimization technique in terms of average CPU time versus $P_{\mathrm{R}}/N_{o}$. The plot reveals that the computational complexity of the optimization algorithm stabilizes after $15$ dB $P_{\mathrm{R}}/N_{o}$ values. This is because, at lower $P_{\mathrm{R}}/N_{o}$, the ISAC system cannot maintain the minimum threshold for all targets and UEs, placing the problem in the infeasible region. Once in the feasible region, the average CPU time fluctuates slightly between $0.02$ and $0.04$ seconds.
@@ VIII Conclusion
In conclusion, this paper has proposed and validated a novel KLD-based framework for optimizing ISAC systems. The two considered optimization techniques, namely, KLD-based radar waveform optimization and KLD-based communication precoding optimization, demonstrated significant improvements over non-optimized scenarios. Radar waveform KLD-based optimization yielded substantial enhancements in both target detection and communication performance, as evidenced by improved ROC curves and reduced BER across all UEs. Communication KLD-based precoding optimization, while primarily benefiting the communication subsystem, also showed modest improvements in radar performance, particularly at lower SNRs. Both techniques exhibited robust performance across varying SNR levels, with radar waveform KLD-based optimization demonstrating stable computational efficiency. Moreover, the trade-off analysis revealed that incorporating the communication signal in radar detection significantly enhances overall ISAC system performance. These findings provide valuable insights for designing and optimizing shared deployment ISAC systems, offering a promising approach to balancing sensing and communication requirements in next-generation wireless networks. The proposed KLD-based framework contributes to the holistic design of ISAC systems, paving the way for more efficient and versatile wireless communications and sensing technologies.
@@ Appendix A Gradient Derivations
This appendix provides detailed derivations for the gradients used in our optimization problem.

### v2

@@ VII-C Communication Waveform KLD-based Optimisation
The communication waveform KLD-based optimisation algorithm is configured similarly to the radar algorithm, with a maximum iteration limit of $\mathrm{max\_iter}=1000$ and a convergence tolerance of $\epsilon=10^{-6}$. The barrier parameter is initialised at $\tilde{\mu_{0}}=1$ and the barrier reduction factor at $\tilde{\gamma}=0.1$. The KLD lower bounds for all targets and UEs, $\underline{A}_{t}$ and $\underline{B}_{k}$, are set to 10 bits.
Fig. 7 illustrates the performance of the radar and communication subsystems using the optimised communication waveform. Fig. 7.a shows the achievable $\mathrm{KLD_{r}}$ for each target as a function of $P_{\mathrm{R}}/N_{o}$ for power allocations of $P_{\mathrm{r}}=\{0.7,0.5,0.3\}$. As $P_{\mathrm{R}}/N_{o}$ increases, $\mathrm{KLD_{r}}$ improves across all targets, with target $\mathrm{T}_{3}$ consistently achieving the highest $\mathrm{KLD_{r}}$, followed by $\mathrm{T}_{2}$ and $\mathrm{T}_{1}$. For example, at $P_{\mathrm{R}}/N_{o}=10$ dB and $P_{\mathrm{r}}=0.7$, the $\{\mathrm{KLD}_{\mathrm{r},1},\mathrm{KLD}_{\mathrm{r},2},\mathrm{KLD}_{\mathrm{r},3}\}$ values are approximately $\{8.42,10.78,12.64\}$, increasing to $\{38.26,42.14,47.53\}$ at $P_{\mathrm{R}}/N_{o}=20$ dB. The linear improvement in $\mathrm{KLD_{r}}$ beyond $15$ dB indicates stable radar performance despite the suboptimal waveform. Fig. 7.b depicts the communication subsystem performance with $\mathrm{KLD_{c}}$ versus $P_{\mathrm{R}}/N_{o}$ for $P_{\mathrm{c}}=\{0.7,0.5,0.3\}$. $\mathrm{UE}_{3}$ achieves the highest $\mathrm{KLD_{c}}$, followed by $\mathrm{UE}_{2}$ and $\mathrm{UE}_{1}$. At $P_{\mathrm{R}}/N_{o}=10$ dB and $P_{\mathrm{c}}=0.7$, the $\{\mathrm{KLD}_{\mathrm{c},1},\mathrm{KLD}_{\mathrm{c},2},\mathrm{KLD}_{\mathrm{c},3}\}$ values are approximately $\{45.87,63.92,86.14\}$, rising to $\{128.43,159.47,182.76\}$ at $P_{\mathrm{R}}/N_{o}=20$ dB. The communication performance shows less sharp improvement at higher $P_{\mathrm{R}}/N_{o}$ due to the need to balance radar subsystem thresholds with communication enhancements across different power allocations.
@@
Fig. 8.a shows the ROC curves for the three targets at different $P_{\mathrm{R}}/N_{o}$ levels. At $P_{\mathrm{R}}/N_{o}=10$ dB and $P_{FA}=10^{-3}$, the $P_{D}$ values are $\{0.397,0.231,0.454\}$ for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$, respectively, compared to $\{0.217,0.109,0.273\}$ for the non-optimal case. These results indicate that communication optimisation enhances detection capabilities at lower $P_{\mathrm{R}}/N_{o}$ compared to non-optimisation, as shown in Fig. 2. Fig. 8.b depicts the BER performance for three UEs versus $P_{\mathrm{R}}/N_{o}$. The BER decreases with increasing $P_{\mathrm{R}}/N_{o}$, reflecting improved communication performance. $\mathrm{UE}_{3}$ consistently achieves the lowest BER, followed by $\mathrm{UE}_{2}$ and $\mathrm{UE}_{1}$. In the non-optimal scenario, $P_{e}$ saturates at $8\times 10^{-3}$ after 15 dB (Fig. 1), while with optimisation, $P_{e}$ drops below $10^{-4}$ by 15 dB across all UEs. Although optimised waveforms provide a performance gain over the non-optimal method, the gain is less pronounced than with optimised radar waveforms due to radar interference.
@@
Fig. 9.a illustrates the trade-off between both subsystems using the optimised communication waveform, showing $\mathrm{KLD_{c}}$ versus $\mathrm{KLD_{r}}$ for different $P_{\mathrm{R}}/N_{o}$ values. As $P_{\mathrm{R}}/N_{o}$ increases from left to right, both $\mathrm{KLD_{c}}$ and $\mathrm{KLD_{r}}$ improve, indicating enhanced performance in both subsystems. This highlights the robustness of the optimisation technique in balancing radar and communication performance. In contrast, the non-optimal case in Fig. 3.b shows $\mathrm{KLD_{c}}$ saturating around 11 bits. Fig. 9.b presents the average CPU time for the communication waveform optimisation technique versus $P_{\mathrm{R}}/N_{o}$. The computational complexity stabilises after $15$ dB as the ISAC system transitions from an infeasible region, where minimum thresholds cannot be met for all targets and UEs, to a feasible region, with average CPU time fluctuating between $0.02$ and $0.04$ seconds.
@@ VIII Conclusion
This paper has proposed and validated a novel KLD-based framework for optimising ISAC systems. The two considered optimisation techniques, namely, KLD-based radar waveform optimisation and KLD-based communication precoding optimisation, demonstrated significant improvements over non-optimised scenarios. Radar waveform KLD-based optimisation yielded substantial enhancements in both target detection and communication performance, as evidenced by the ROC curves and reduced BER across all UEs. Communication KLD-based precoding optimisation, while primarily benefiting the communication subsystem, also showed modest improvements in radar performance, particularly at lower SNRs. Both techniques exhibited robust performance across varying SNR levels, with radar waveform KLD-based optimisation demonstrating stable computational efficiency. Moreover, trade-off analysis revealed that incorporating the communication signal in radar detection significantly enhances the overall ISAC system performance. These findings provide valuable insights for designing and optimising shared deployment ISAC systems, offering a promising approach to balance sensing and communication requirements in future wireless networks. The proposed framework provides the means for a holistic ISAC design, paving the way for more efficient and versatile wireless communications and sensing technologies.
@@ Appendix A Gradient Derivations
This appendix provides detailed derivations for the gradients used in our optimisation problem.